OpenStudy (anonymous):

Division Properties of Exponents If I have 8 on the top with y^7 z^6 and 4 with y^6z^5 would I divide the whole number too?

OpenStudy (campbell_st):

use the system, 1. Numbers 2. Letters 3. Powers 4. Signs so for the numbers its 8/4 = 2 letters y^7/y^6 = y z^6/z^5 = z powers... both 1 signs everything is positive now combine for the answer...

OpenStudy (anonymous):

so 2yz?

OpenStudy (campbell_st):

thats it

OpenStudy (anonymous):

Thank you!